Please help me I need to know how to do it as well
Jane Corporation produces model toy cars Each sells for $31.05. Its variable cost per unit is $11.38 Assuming it has a fixed cost of $639 275. What is the breakeven point for Jane Corporation? units
Answer #1

Break even point = Fixed cost/contribution margin per unit

Contribution margin per unit = 31.05-11.38 = 19.67 per unit

Break even point = 639275/19.67 = 32500 units